The nominal price of a college education rose to $15,212 in 2009. The CPI was 107.6 in 1985, 130.7 in 1990, and 144.0 in 2009.
A)What was the real price of a college education in 2009 in terms of 1985 dollars?
B) What was the real price of a college education in 2009 In terms of 1990 dollars?

MUST SHOW ALL OF YOUR WORK
1
Expert's answer
2019-07-30T09:20:41-0400

Real Price"_{z}" =  (Nominal Price"_{z}" ) ×  "\frac{CPI _{base year} }{CPI{} _{z} }"

A. Real price in terms of 1985 dollar = 15212× ("\frac{144.0}{130.7}")= $20358.07

B. Real price in terms of 1990 dollar =15212 × ("\frac{144.0}{130.7}")= $16759.97